OpenStudy (anonymous):

Write an equation of the line with the given slope that contains the given point.. (3,-6) slope 1/3

OpenStudy (nikato):

Do u know the point slope form? U can just plug in those numbers into that equation

OpenStudy (anonymous):

is it y=mx+b?

OpenStudy (nikato):

No that's slope intercept. But u can also use that too

OpenStudy (the_fizicx99):

No, thats to find your slope M. y-y1 = m(x-x1)

OpenStudy (anonymous):

Can I go 1/3=3+-6?

OpenStudy (nikato):

No, r u trying to plug it into ur equation y=mx+b?

OpenStudy (the_fizicx99):

Nope. y+6=1/3(x-3)

OpenStudy (nikato):

If u r, it's acutaully -6=1/3(3)+b And then u would solve for b

OpenStudy (anonymous):

how did u do that? & would that be the only yhing i do?

OpenStudy (the_fizicx99):

No @nikato

OpenStudy (nikato):

There's nothing wrong @tHe_FiZiCx99 I'm just using the equation @rosagonzalez wanted to use

OpenStudy (the_fizicx99):

Why would you use y=mx+b...

OpenStudy (nikato):

Why can't u?

OpenStudy (the_fizicx99):

Because she has her slope and points. ask @shamil98

OpenStudy (nikato):

Sooooo. U can still use it. It'll just be one step longer

OpenStudy (the_fizicx99):

Ok, you can explain with your method. Gl :p

OpenStudy (anonymous):

I want to use the easiest way. I have a test tomrrow morning. .-.

OpenStudy (nikato):

@tHe_FiZiCx99 his way would be faster if ur teacher wants it in point slope form. It depends what equation ur teacher wants ur answer in

OpenStudy (anonymous):

ok. thank you. i'll try using both and see which one is easy for me

OpenStudy (anonymous):

so lets say it says; write an equation in slope-intercept form of the line though point (-10,1) with slope -5. would i go y-1=-5(x-10) then it would be y-1=-5x-50.. so the answer would be y=-5x-49

OpenStudy (anonymous):

@nikato @tHe_FiZiCx99

OpenStudy (shamil98):

(3,-6) slope 1/3 Just plug in your given values into \[\huge y-y_1=m(x-x_1)\] and simplify :)

OpenStudy (shamil98):

And to confirm your other question (-10,1) , m = -5 \[\huge y - 1 = -5(x-(-10))\] \[\huge y - 1 = -5x - 50\] \[\huge y = -5x - 49\] Correct! Good job.

OpenStudy (anonymous):

thank you so much!(:

OpenStudy (the_fizicx99):

Seee!!

OpenStudy (nikato):

What? Both ways work

OpenStudy (the_fizicx99):

Mine was faster is what I was trying to prove :p

OpenStudy (nikato):

I never said mines was faster